Calculate the pH of a mixture containing 50ml of 1(N) HCl and 30 ml 1(N) NaOH .((Ans.. 0.6)).
 

Grade:11

2 Answers

Rohan Gujar
5 years ago
Just use formula [H+]=(M1V1+M2V2)/(V1+V2) 
So, [H+]=(50-30)/80.    [N=M×n-factor]
[H+]=0.25.      So pH= -log[H+]
                                   = -log(0.25)=0.6 Ans.
Arun
5 years ago
Dear student
 

Mole = concentration x volume

mole of NaOH = 30ml x 1 = 30.0

mole of HCl = 50ml x 1= 50.0

30 mole of NaOH exactly neutralized by 30 mole of HCl.

NaOH + HCl = NaCl + H2O

that is 50 -30 = 20 HCl mole

Con of H+ = no of mole of H+ / total volume of solution (50+ 30 ml = 80 ml)

= 20/80 = 0.25

pH = - log ( H+)

= - log (1/4) = 0.6021

s0 the pH of solution will be 0.6021

 

 

Regards

Arun